find the solution to the exponential equation correct to four decimal equations.
5+3^(5x)=13

Answer by Fombitz(13823)

5+3^(5x)=13
3^(5x)=8
5x=log(8,3)
x=(1/5)log(8,3)
log(8,3) is log 8 base 3.
log(8,3)=log(8)/log(3) in base 10.
x=(1/5)(log(8)/log(3))
or approximately
x=0.379